Four point charges are equally separated by 1.00 m in air. If q1=3.0nC, q2=2.0nC, and q4=-5.0nC. What is the electric field at a point halfway between q2 and q3?


1
Expert's answer
2022-02-09T06:28:59-0500

The electric field at a point halfway between q2 and q3 is


"E=\dfrac1{4\pi\epsilon_0r^2}\bigg(\dfrac{q_1}{1.5^2}+\dfrac{q_2}{0.5^2}+\dfrac{q_3}{0.5^2}+\dfrac{q_4}{1.5^2}\bigg),\\\space\\
E=(103.9+3.595·10^{10}·q_3)\text{ V/m}."